OpenStudy (anonymous):

Solve for x: 2 over 5 plus 3 over 5 x equals x plus 5 over 10

OpenStudy (anonymous):

\[\frac{2}{5}+\frac{3}{5}x=\frac{x+5}{10}\]?

OpenStudy (lgbasallote):

\[\huge \frac{2}{5 + \frac{3}{5x}} = \frac{x+5}{10}\] i think

OpenStudy (anonymous):

i doubt it

OpenStudy (anonymous):

its the first one but the x is next to the 5 like the second

OpenStudy (lgbasallote):

a guy can dream cant he :(

OpenStudy (valpey):

So\[\frac{2}{5}+\frac{3}{5x}=\frac{x+5}{10}\] I'd start with multiplying the whole problem by the LCD of 10x.

OpenStudy (valpey):

The three denominators are 5, 5x, and 10. The least common multiple of these three things is 10x. When you multiply each fraction by 10x you will be left with no fractions.

OpenStudy (anonymous):

oh so they cancel out and its no solution?

OpenStudy (valpey):

No. You will have to do some cancelling, but you will still have some terms left.